problems based on laws of chemical conservation

Grade:9

1 Answers

Gaurav
9 years ago
Sameple question for law of conservation of mass

Quesion .
5.2 g of CaCo hen heated produced 1.99 g of Carbon dioxide and the residue (CaO) left behind weighs 3.2g. Show that these results illustrate the law of conservation of mass.

Solution:
Weight ofCaCo taken = 5.2 g

Total weight of the products (CaO +CO )= 3.20+ 1.99 = 5.19 g

Difference between the wt. of the reactant and the total wt. of the products= 5.20 – 5.19 =0.01 g.

This small difference may be due to experimental error.

Thus law of conservation of mass holds good within experimental errors.